# SemiEvol: Semi-supervised Fine-tuning for LLM Adaptation

### v1

@@ Abstract
Supervised fine-tuning (SFT) is crucial in adapting large language models (LLMs) to a specific domain or task. However, only a limited amount of labeled data is available in practical applications, which poses a severe challenge for SFT in yielding satisfactory results. Therefore, a data-efficient framework that can fully exploit labeled and unlabeled data for LLM fine-tuning is highly anticipated. Towards this end, we introduce a semi-supervised fine-tuning framework named SemiEvol for LLM adaptation from a propagate-and-select manner. For knowledge propagation, SemiEvol adopts a bi-level approach, propagating knowledge from labeled data to unlabeled data through both in-weight and in-context methods. For knowledge selection, SemiEvol incorporates a collaborative learning mechanism, selecting higher-quality pseudo-response samples. We conducted experiments using GPT-4o-mini and Llama-3.1 on seven general or domain-specific datasets, demonstrating significant improvements in model performance on target data. Furthermore, we compared SemiEvol with SFT and self-evolution methods, highlighting its practicality in hybrid data scenarios. [^1] GitHub repository: https://github.com/luo-junyu/SemiEvol.
SemiEvol: Semi-supervised Fine-tuning for LLM Adaptation
@@ 1 Introduction
In practice, however, it often presents a hybrid situation, where a small amount of labeled data coexists with a relatively larger volume of unlabeled data. On the one hand, when deploying LLMs to new target tasks, a limited amount of task-specific annotations can be valuable without incurring excessive costs Perlitz et al. (2023); Kung et al. (2023). On the other hand, during the continuous inference process of LLMs, a substantial amount of unlabeled data accumulates Tao et al. (2024); Honovich et al. (2023); Wang et al. (2023b). Effectively leveraging the labeled data to enhance model performance on unlabeled data, while simultaneously selecting high-quality unlabeled samples, can improve LLMs’ performance in target scenarios, offering substantial practical utility. Therefore, we aim to address the following question:
@@
  • To the best of our knowledge, we are the first to study a practical problem of semi-supervised fine-tuning, aiming to adapt LLMs into different domains data-efficiently.
@@ 2.1 Supervised Fine-tuning
2.2 Background and Problem Definition: Semi-supervised Fine-tuning
@@ 3.3 Collaborative Learning
To further exploit unlabeled data, we designed a collaborative learning framework tailored for LLMs. This framework utilizes the inherent capabilities of LLMs for self-justify to obtain high-confidence pseudo-responses from unlabeled data. Some concurrent works also attempt to use LLMs for similar functionality Wang et al. (2024a), while their focus differs from ours.
@@ 3.4 Knowledge Adaptive Selection
Finally, we use the high-quality pseudo-responses to fine-tune the model, which can enhance its performance and adaptability on the target task:
${\mathcal{M}}_{\text{evol}}=FT\left({\mathcal{M}}_{\text{warm}},{\mathcal{D}}_{\text{selected}}\right)\,,$ (9)
where ${\mathcal{M}}_{\text{warm}}$ is the model obtained after initial fine-tuning in Eq. 1, and $FT$ denotes the fine-tuning process.
By focusing on these high-quality assessed data, we enhance the model’s performance and adaptability on the target task while reducing the influence of noisy or erroneous information.
@@ 4.1.1 Datasets
We employed both general-purpose and domain-specific evaluation datasets to provide a comprehensive assessment. These datasets encompass a variety of tasks, including multiple-choice questions, reasoning, numerical computations, etc.. Specifically, our general evaluation datasets include MMLU Hendrycks et al. (2020), MMLU-Pro Wang et al. (2024c), and ARC Clark et al. (2018), while domain-specific datasets comprise FPB Malo et al. (2014), USMLE Jin et al. (2021), PubMedQA Jin et al. (2019), and ConvFinQA Chen et al. (2022), covering various fields such as finance and healthcare. This diverse selection enables a thorough evaluation of the model’s performance across different task types and knowledge domains.
@@ 4.1.2 Backbones and Baselines
Baselines. We evaluated our method against baselines from several categories: (1) Vanilla, which involves testing solely through API calls or using the original model; (2) Supervised Fine-tuning (SFT) Hu et al. (2021); Wei et al. (2021), which adapts the model to the target task using the labeled data; (3) Self-Evolution Methods (SelfEvol), which enhance LLM capabilities using additional unlabeled data. We compare with Reflection-Llama Li et al. (2024) [^2] https://huggingface.co/Solshine/reflection-llama-3.1-8B and Hermes-3 Teknium et al. (2024) [^3] https://huggingface.co/NousResearch/Hermes-3-Llama-3.1-8B, both of which evolve from the Llama-3.1-8B model; (4) Domain Adaptation Methods, including AdaptLLM Cheng et al. (2024b) and InstructPT Cheng et al. (2024a), utilize domain-specific data (e.g., finance and medical). We select models adapted to corresponding domains for testing, all with comparable parameter counts of 8B; (5) Inference-time enhancement methods, such as Retrieval Augmented Generation (RAG) Lewis et al. (2020), including BM25 Jones et al. (2000) and FAISS Douze et al. (2024) algorithms. We also compare with MemoryLLM Wang et al. (2024b), with the nearest labeled sample as memory;
@@ 4.1.3 Implementation Details
For the setting of semi-supervised fine-tuning of LLMs, we have ${\mathcal{D}}_{\text{labeled}}$, ${\mathcal{D}}_{\text{unlabeled}}$ and ${\mathcal{D}}_{\text{test}}$. The data proportion in our experiments is $labeled:unlabeled:test=2:6:2$ and will be further discussed in Section 4.3.6. The answer information for ${\mathcal{D}}_{\text{unlabeled}}$ is inaccessible in our setting. We fine-tuned Llama-3.1-8B using Low-Rank Adaptation (LoRA) Hu et al. (2021) and applied fine-tuning with the official API for GPT-4o-mini [^4] https://platform.openai.com/finetune.. All fine-tuning processes take $2$ epochs. $n$ is set to $4$ and $\theta$ is set to $50\%$, with further investigation planned in subsequent experiments. Detailed training configurations are provided in the Appendix.
@@ 4.3.7 Discussion on Iterative Evolution
We verify the model’s iterative evolution capability, as illustrated in Figure 7. After applying SemiEvol, we utilized the labeled data and pseudo-response data as new labeled data, initiating a fresh round of SemiEvol on the previously filtered unlabeled data. By the fourth iteration, we had utilized $94.75\%$ of the unlabeled data, resulting in further performance improvements in the target scenario. The model’s performance on MMLU-Pro exceeded $55\%$. This iterative evolution capability further demonstrates the practicality of SemiEvol.
@@ 5.1 Data Engineering for SFT
With the rapid advancement of Large Language Models (LLMs) Zhao et al. (2023), researchers have discovered that employing suitable data for Supervised Fine-Tuning (SFT) can enhance model performance on downstream tasks Taori et al. (2023); Longpre et al. (2023). Some researchers focus on data selection Bhatt et al. (2024); Zhou et al. (2024); Xia et al. (2024); Bukharin and Zhao (2023), aiming to improve data quality to boost model effectiveness within limited training budgets. Others concentrate on data synthesis Xu et al. (2023); Mukherjee et al. (2023); Chung et al. (2024); Honovich et al. (2022); Cheng et al. (2023), attempting to enhance models’ instruction-following capabilities through synthesized instruction data. Complementary to these approaches, SemiEvol focuses on LLMs’ ability to continuously evolve in real-world scenarios, relying solely on their inherent capabilities. It effectively utilizes small amounts of labeled data to improve model evolution performance.
@@ 5.2 Semi-supervised Learning
Semi-supervised learning aims to reduce the annotation cost during model training Zhu (2005); Tarvainen and Valpola (2017), which has received increasing attention in various fields such as text classification Duarte and Berton (2023); Thangaraj and Sivakami (2018); Linmei et al. (2019) and neural machine translation Cheng et al. (2016); Pham et al. (2023). Current semi-supervised learning approaches can be mainly divided into two types, i.e., pseudo-labeling Lee et al. (2013) and consistency regularization Sohn et al. (2020); Berthelot et al. (2019). Pseudo-labeling approaches usually add extra unlabeled data into the labeled dataset by leveraging the labels predicted by the model. Recent studies attempt different techniques to enhance pseudo-labeling such as considering adaptive thresholds Zhang et al. (2024); Rhee and Cho (2019) and class imbalance Wang et al. (2023a). In contrast, consistency regularization aims to encourage the consistency of predictions under different perturbations. However, these approaches focus on classification problems Shi et al. (2023), which cannot be applied to LLM fine-tuning. To tackle this issue, we propose a new framework SemiEvol in a propagate-and-select manner for LLM adaptation.
@@ 6 Conclusion
Ethics Statement
Our research adheres to the ACL Code of Ethics. All datasets and language models used in this study are publicly available. The code and related materials will be appropriately released to ensure transparency and reproducibility of our work.

### v2

# Semi-supervised Fine-tuning for Large Language Models

@@ Abstract
Supervised fine-tuning (SFT) is crucial in adapting large language models (LLMs) to a specific domain or task. However, only a limited amount of labeled data is available in practical applications, which poses a severe challenge for SFT in yielding satisfactory results. Therefore, a data-efficient framework that can fully exploit labeled and unlabeled data for LLM fine-tuning is highly anticipated. Towards this end, we introduce a semi-supervised fine-tuning (SemiFT) task and a framework named SemiEvol for LLM alignment from a propagate-and-select manner. For knowledge propagation, SemiEvol adopts a bi-level approach, propagating knowledge from labeled data to unlabeled data through both in-weight and in-context methods. For knowledge selection, SemiEvol incorporates a collaborative learning mechanism, selecting higher-quality pseudo-response samples. We conducted experiments using GPT-4o-mini and Llama-3.1 on seven general or domain-specific datasets, demonstrating significant improvements in model performance on target data. Furthermore, we compared SemiEvol with SFT and self-evolution methods, highlighting its practicality in hybrid data scenarios.
Semi-supervised Fine-tuning for Large Language Models
@@ 1 Introduction
In practice, however, it often presents a hybrid situation, where a small amount of labeled data coexists with a relatively larger volume of unlabeled data. On the one hand, when deploying LLMs to new target tasks, a limited amount of task-specific annotations can be valuable without incurring excessive costs Perlitz et al. (2023); Kung et al. (2023). On the other hand, during the continuous inference process of LLMs, a substantial amount of unlabeled data accumulates Tao et al. (2024); Honovich et al. (2023); Wang et al. (2023). Effectively leveraging the labeled data to enhance model performance on unlabeled data, while simultaneously selecting high-quality unlabeled samples, can improve LLMs’ performance in target scenarios, offering substantial practical utility. Therefore, we aim to address the following question:
@@
  • To the best of our knowledge, we are the first to study a practical problem of semi-supervised fine-tuning (SemiFT), aiming to adapt LLMs into different domains data-efficiently.
@@ 2.1 Supervised Fine-tuning
2.2 Background and Problem Definition: Semi-supervised Fine-tuning (SemiFT)
@@ 3.3 Collaborative Learning
To further exploit unlabeled data, we designed a collaborative learning framework tailored for LLMs. This framework utilizes the inherent capabilities of LLMs for self-justify to obtain high-confidence pseudo-responses from unlabeled data. Some concurrent works also attempt to use LLMs for similar functionality Wang et al. (2024b), while their focus differs from ours.
@@ 3.4 Knowledge Adaptive Selection
Finally, we combine the selected pseudo-labeled data with the original labeled data to fine-tune the base model, which can enhance its performance and adaptability on the target task:
${\mathcal{M}}_{\text{evol}}=FT\left({\mathcal{M}}_{\text{base}},{\mathcal{D}}_{\text{selected}}~{}\cup~{}{\mathcal{D}}_{\text{labeled}}\right)\,,$ (9)
where ${\mathcal{M}}_{\text{base}}$ is the pre-trained LLM, and $FT$ denotes the fine-tuning process.
By leveraging both high-quality pseudo-labeled data and original labeled data, we enhance the model’s performance and adaptability on the target task while reducing the influence of noisy or erroneous information.
@@ 4.1.1 Datasets
We employed both general-purpose and domain-specific evaluation datasets to provide a comprehensive assessment. These datasets encompass a variety of tasks, including multiple-choice questions, reasoning, numerical computations, etc.. Specifically, our general evaluation datasets include MMLU Hendrycks et al. (2020), MMLU-Pro Wang et al. (2024d), and ARC Clark et al. (2018), while domain-specific datasets comprise FPB Malo et al. (2014), USMLE Jin et al. (2021), PubMedQA Jin et al. (2019); Jiang et al. (2024a), and ConvFinQA Chen et al. (2022), covering various fields such as finance and healthcare. This diverse selection enables a thorough evaluation of the model’s performance across different task types and knowledge domains.
@@ 4.1.2 Backbones and Baselines
Baselines. We evaluated our method against baselines from several categories: (1) Vanilla, which involves testing solely through API calls or using the original model; (2) Supervised Fine-tuning (SFT) Hu et al. (2021); Wei et al. (2021), which adapts the model to the target task using the labeled data; (3) Self-Evolution Methods (SelfEvol), which enhance LLM capabilities using additional unlabeled data. We compare with Reflection-Llama Li et al. (2024) [^1] https://huggingface.co/Solshine/reflection-llama-3.1-8B and Hermes-3 Teknium et al. (2024) [^2] https://huggingface.co/NousResearch/Hermes-3-Llama-3.1-8B, both of which evolve from the Llama-3.1-8B model; (4) Domain Adaptation Methods, including AdaptLLM Cheng et al. (2024b) and InstructPT Cheng et al. (2024a), utilize domain-specific data (e.g., finance and medical). We select models adapted to corresponding domains for testing, all with comparable parameter counts of 8B; (5) Inference-time enhancement methods, such as Retrieval Augmented Generation (RAG) Lewis et al. (2020), including BM25 Jones et al. (2000) and FAISS Douze et al. (2024) algorithms. We also compare with MemoryLLM Wang et al. (2024c), with the nearest labeled sample as memory;
@@ 4.1.3 Implementation Details
For the setting of semi-supervised fine-tuning of LLMs, we have ${\mathcal{D}}_{\text{labeled}}$, ${\mathcal{D}}_{\text{unlabeled}}$ and ${\mathcal{D}}_{\text{test}}$. The data proportion in our experiments is $labeled:unlabeled:test=2:6:2$ and will be further discussed in Section 4.3.6. The answer information for ${\mathcal{D}}_{\text{unlabeled}}$ is inaccessible in our setting. We fine-tuned Llama-3.1-8B using Low-Rank Adaptation (LoRA) Hu et al. (2021) and applied fine-tuning with the official API for GPT-4o-mini [^3] https://platform.openai.com/finetune.. All fine-tuning processes take $2$ epochs. $n$ is set to $4$ and $\theta$ is set to $50\%$, with further investigation planned in subsequent experiments. Our dataset is publicly available at Hugging Face [^4] https://huggingface.co/datasets/luojunyu/SemiEvol.
@@ 4.3.7 Discussion on Iterative Evolution
We verify the model’s iterative evolution capability, as illustrated in Figure 7. After applying SemiEvol, we utilized the labeled data and pseudo-response data as new labeled data, initiating a fresh round of SemiEvol on the previously filtered unlabeled data. By the fourth iteration, we had utilized most of the unlabeled data, resulting in further performance improvements in the target scenario. This iterative evolution capability further demonstrates the practicality of SemiEvol.
@@ 5.1 Data Engineering for SFT
With the rapid advancement of Large Language Models (LLMs) Zhao et al. (2023), researchers have discovered that employing suitable data for Supervised Fine-Tuning (SFT) can enhance model performance on downstream tasks Taori et al. (2023); Longpre et al. (2023); Hou et al. (2024); Luo et al. (2024b); Jiang et al. (2024b). Some researchers focus on data selection Bhatt et al. (2024); Parkar et al. (2024); Xia et al. (2024); Bukharin and Zhao (2023), aiming to improve data quality to boost model effectiveness within limited training budgets. Others concentrate on data synthesis Mukherjee et al. (2023); Chung et al. (2024); Honovich et al. (2022); Cheng et al. (2023), attempting to enhance models’ instruction-following capabilities through synthesized instruction data. Researchers also shifted their focus to model self-evolution Tao et al. (2024); Madsen et al. (2024). These include self-instruction Wang et al. (2022) and self-play Chen et al. (2024) techniques that enable models to acquire task-specific capabilities without extensive annotated data. Complementary to these approaches, SemiEvol focuses on LLMs’ ability to continuously evolve in real-world semi-supervised fine-tuning (SemiFT) scenarios, relying solely on their inherent capabilities. It effectively utilizes small amounts of labeled data to improve model evolution performance.
@@ 5.2 Semi-supervised Learning
Semi-supervised learning aims to reduce the annotation cost during model training Zhu (2005); Tarvainen and Valpola (2017); Ju et al. (2024); Yang et al. (2024); Feng et al. (2024), which has received increasing attention in various fields such as text classification Duarte and Berton (2023); Thangaraj and Sivakami (2018); Linmei et al. (2019) and neural machine translation Cheng et al. (2016); Pham et al. (2023). Current semi-supervised learning approaches can be mainly divided into two types, i.e., pseudo-labeling Lee et al. (2013) and consistency regularization Sohn et al. (2020); Berthelot et al. (2019). Pseudo-labeling approaches usually add extra unlabeled data into the labeled dataset by leveraging the labels predicted by the model. Recent studies attempt different techniques to enhance pseudo-labeling such as considering adaptive thresholds Zhang et al. (2024); Rhee and Cho (2019) and class imbalance Luo et al. (2024a); Wang et al. (2024a). In contrast, consistency regularization aims to encourage the consistency of predictions under different perturbations. However, these approaches focus on classification problems Shi et al. (2023), which cannot be applied to LLM fine-tuning. To tackle this issue, we propose a new framework SemiEvol in a propagate-and-select manner for LLM adaptation.
@@ 6 Conclusion
Acknowledgement
This paper is partially supported by the National Key Research and Development Program of China with Grant No. 2023YFC3341203 as well as the National Natural Science Foundation of China with Grant Numbers 62276002 and 62306014.
Appendix A Algorithm
In Algorithm 1, we present the comprehensive algorithmic process of SemiEvol, which incorporates a bi-level framework for knowledge propagation and selection. This process ultimately yields the evolved model, ${\mathcal{M}}_{evol}$.
Algorithm 1 Algorithm of SemiEvol
Require: Labeled data ${\mathcal{D}}_{labeled}$, Unlabeled data ${\mathcal{D}}_{unlabeled}$, LLM ${\mathcal{M}}$;
Ensure: Evolved LLM ${\mathcal{M}}_{evol}$;
1:  // In-Weight Knowledge Propagation
2:  Fine-tune ${\mathcal{M}}$ on ${\mathcal{D}}_{labeled}$, obtain ${\mathcal{M}}_{warm}$;
3:  // Collaborative Learning
4:  for m = 1, $\cdots$, $n$ do
5:     // In-Context Propagation
6:     Get the prediction $\left\{y^{m}_{j}\right\}$ as Eq. 4;
7:  end for
8:  Self-Justify to generate pseudo-responses;
9:  // Adaptive Selection
10:  Select pseudo-responses with entropy as Eq. 8, obtain ${\mathcal{D}}_{selected}$;
11:  Fine-tune on ${\mathcal{D}}_{selected}$, obtain ${\mathcal{M}}_{evol}$;
Appendix B Experimental Settings
We will detail the experimental process, including parameter settings, prompt configurations, and computational resource consumption.
B.1 Parameter Settings
Inference Process For commercial methods (e.g., GPT4o-mini), we utilize API calls. For open-source models (e.g., Llama3.1 8B), we employ the vLLM Kwon et al. (2023) framework locally for inference. The inference parameters are as Table 4.
Parameter Value
Model Name Llama3.1 8B Instruct
Computing Arch. NVIDIA A100-80GB
Max Sequence Length 4096
Table 4: Parameters configuration during inference.
In the collaborative learning process, we employ $n$ LLMs with diverse configurations for mutual learning, setting their temperature to $1$. These configurations are sampled from the following options: (1) Utilization of ${\mathcal{M}}_{w}arm$: $\{0,1\}$; (2) Number of labeled samples referenced for in-context knowledge propagation: $\{0,1,2,3\}$.
Fine-tuning Process. We use platform APIs for commercial methods fine-tuning and LlamaFactory Zheng et al. (2024) for open-source models fine-tuning. Both of them are fine-tuned for $2$ epochs. Commercial models use default settings (adaptively configured by OpenAI based on the task). For open-source models, our hyperparameter settings are as Table 5, all of which are the default parameters from LlamaFactory.
Parameter Value
Model Name Llama3.1 8B Instruct
Computing Arch. NVIDIA A100-80GB
Fine-tuning Type LoRA
Gradient Accumulation Steps 8
Learning Rate 1e-4
$\theta$ in Adaptive Selection $50\%$
Table 5: Hyperparameter settings for fine-tuning open-source models.
B.2 Instruction Settings
We employ concise instructions for inference as shown in Table 6. During this process, we present questions to the LLM and elicit responses.
Instruction Template
Multiple Choice
Answer the multiple-choice question.
Your response should be in the format: ’Answer: LETTER’ (without quotes).
Question:
{question}
Options:
{options}
Free-from QA
Answer the following question.
Output the value in the format: ’Answer: VALUE’ (without quotes).
Question:
{question}
Options:
{options}
[…]
Table 6: Instruction templates for different types of questions.
We extract answers using regular expression matching. For character-based answers, we check for exact matches. For numerical answers, we assess whether they fall within an acceptable error margin (maximum error of 1e-2).
During the SemiEvol process, we require additional instructions for tasks such as self-justification and in-context knowledge propagation. For these supplementary commands, we provide instruction templates in a table for reference.
Instruction Template
Self-Justify
Here are the multiple answers to the question.
Please consider them thoroughly and give me the correct answer. Your response should be in the following format:
’Answer: LETTER’ (without quotes).
Question:
{question}
Options: {options}
Multiple Answers:
{answers}
Now, please give me the final correct answer:
In-context Knowledge Propagation
You are an expert in the question answering. Below are some examples of questions and their corresponding answer.
{reference}
Table 7: Instruction templates for SemiEvol process.